# Contextualize-then-Aggregate: Circuits for In-Context Learning in Gemma-2 2B

### v1

@@ 3.2 Which Information is Routed?
For each edge, we specifically contrast the following hypotheses: An edge from some fewshot example $(x_{i},y_{i})$ might transport causally relevant information about (1) the specific tokens, (2) the functional relationship between $x_{i}$ and $y_{i}$ (but not their token identities), (3) the type of the $x_{i}$ or $y_{i}$ tokens without regard to their functional relationship.
@@
We vary the information about the input that is available to a particular edge by constructing a contrastive input sample on which certain aspects of the information have been altered (Figure 22), and patching the K and V activations directly feeding into the edge with the corresponding activations computed on that contrastive input (Figure 28). For instance, when patching an edge $y_{i}\rightarrow t_{N+1}$, we modify the key and value activations associated with the positions of $y_{i}$ across all attention heads in the model. These modifications occur specifically when the activations are queried from the positions corresponding to $t_{N+1}$. This patching is done in parallel in each layer and head. All patching is done within the overall circuit identified for the task, i.e., edges not present in the circuit are ablated, and paths not present in the overall circuit play no role even in the patched version. For instance, to determine whether the edges $y_{i}\rightarrow t_{N+1}$ transport contextualized information about preceding fewshots $(x_{i^{\prime}},y_{i^{\prime}})$ ($i^{\prime}<i$), we patch all inputs to this edge at position $y_{i}$ with activations computed on an input where the preceding fewshots have been corrupted; by varying the degree of corruption, we analyze which specific kinds of information are transported (Figure 22). We note that this methodology allows us to focus on causally relevant information used by downstream components. It is thus fundamentally different from supervised probing, which does not distinguish between information used or unused by downstream components. We focused on the 2B model for tractability.
Parallel Subcircuit
@@
We first investigated the information transported by the edges in the Parallel subcircuit, consisting of the $x_{i}\rightarrow y_{i}$, $x_{N+1}\rightarrow t_{N+1}$ and $y_{i}\rightarrow t_{N+1}$ edges. Prior work (Cho et al., 2025; Kharlapenko et al., 2025; Wang et al., 2023) suggests that the $y_{i}\rightarrow t_{N+1}$ edges aggregate task information from the examples. Indeed, when we patched these with arbitrary $x_{i}$ or $y_{i}$ (which disrupts the functional relationship), functionality was largely destroyed (see Figures 3 and 8 for all patching results). In contrast, when we patched the $y_{i}\rightarrow t_{N}$ edges with other fewshots that are valid for the task, functionality was preserved. We also verified that the $x_{i}\rightarrow y_{i}$ edges transport $x_{i}$: patching these edges with corrupted $x_{i}$ leads to total failure. This is to be expected, as in the simple Parallel subcircuit, these edges are the only way for $x_{i}$ to influence the output. These results confirm prior proposals about the aggregation step (Cho et al., 2025; Kharlapenko et al., 2025; Wang et al., 2023) with rigorous causal analysis.
@@
Our methodology allowed us to obtain two further findings. First, the edges causally transport only task information and no token information. Namely, even in the presence of contextualizing edges, i.e., even though there are multiple paths from fewshots to the $t_{N+1}$, performance did not deteriorate even when the $y_{i}\rightarrow t_{N+1}$ and contextualizing edges transported information from different prompts when they were valid for the task. [^6] We emphasize that this is a claim about information that has a causal downstream effect in the circuit, not about information that can be obtained via probing, which is likely to be richer and include token information.
@@
In order to understand the role of these components further, we investigated the functional behavior of the model with ablations applied. Overall, we observed three fallback tasks which accounted for most interpretable errors: copying the query, producing from the correct output space, copying another token from the prompt. With full ablation of the $y_{i}\rightarrow t_{N}$ edges, behavior was often accounted for by copying the query or copying another token from the corrupted input prompt (Figures 14, 16, 15, 18, 17). We note that as the corrupted prompts maintain the separators, the last separator position can still infer the presence of repeated structure, potentially explaining the presence of copying behavior. When patching the edge while preserving the output space, behavior often amounted to reproduction from the correct output space when there is well-delineated output space (e.g., Capitalization, Country-Capital, Person-Sport), or copying from the corrupted prompt when there is none (Copying). This is expected, as such prompts maintain repeated structure in terms of a specific output space, though with disrupted functional relationship.
@@
We next moved to analyzing the information transported by the contextualizing edges, starting with the edges connecting $y$’s ($y_{i}\rightarrow y_{i+1}$ or general $y_{i}\rightarrow y_{j}$). We found that these edges transport both information about $f$ and about $\mathcal{Y}$. Patching the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital only decreases accuracy significantly when $y$ leaves $\mathcal{Y}$, but not under other corruptions (Figure 3). In Capitalization, at 3 shots, replacing with $y\in\mathcal{Y}$ hurts much less than a general ablation of $y$ (accuracy 81% vs 56 %; Figure 8), which is as harmful as full corruption of these edges and leads to a large increase in copying behavior; nonetheless, accuracy still drops compared to ablations keeping $(x,y)\in f$. At 10 shots, Capitalization only responds to ablations where $y$ leaves $\mathcal{Y}$. In both cases, ablating the output space information leads to an increased number of copy-type mistakes (Figure 19).
@@
The $y_{i}\rightarrow y_{j}$ / $y_{i}\rightarrow y_{i+1}$ edges contextualize the representation of each fewshot example $y_{i}$ by transporting the type of $\mathcal{Y}$, and sometimes the task $f$, from $(x_{j},y_{j})$ ($j<i)$.
We next studied the edges connecting inputs $x$’s. Ablating these on Present-Past and Person-Sport turns out to not lead to a statistically significant drop, but the drop is substantive in Country-Capital (Figure 3). Here, ablations only hurt when $x$ left $\mathcal{X}$, showing that $x_{i}\rightarrow x_{j}$ transports the input space, but not (causally) the token $x_{i}$. Such ablations again lead to a substantial fraction of copying responses, i.e., the functional input-output behavior is impacted (Figure 20). Patching prior $x_{i}$’s for these edges, we found that they affect the final prediction both via $x_{i}\rightarrow x_{i+1}\rightarrow y_{i+1}\rightarrow t_{N+1}$ and via $x_{N}\rightarrow x_{N+1}\rightarrow t_{N+1}$ edges (Table 8). Overall:
The $x_{i}\rightarrow x_{i+1}$ edges contextualize the representations of $x_{N+1}$ and of individual fewshot outputs $y_{i}$ with information about $\mathcal{X}$.
@@ 3.3 A Contextualization subcircuit is important in the presence of ambiguity
We conducted a controlled experiment aiming to stress-test the aggregation step. We designed the following hard version of the Present-Past task. We take advantage of the fact that some English verbs have identical present and past tense forms (e.g., “put”, “spread”). In a prompt for the Present-Past task, examples using such verbs are thus ambiguous between the Present-Past task and a simpler Copying task (Figure 23.1). In this experiment, most individual examples do not give full task information, and may in fact provide misleading information. We hypothesize that direct aggregation of uncontextualized per-fewshot representations as hypothesized by prior work may be less useful in this case, because individual examples might, on their own, provide conflicting task information. The contextualization step might help resolve this ambiguity before the aggregation step.
Even when 7 out of 10 fewshot examples are ambiguous, the full model continues to perform the task at very high accuracy (95%). Whereas the aggregation (Parallel)-only circuit had performed at 93% on the standard Present-Past task (and similarly on the Copying task), its performance on the ambiguous prompts dropped to 56% (Table 9).
Figure 5: Two hypotheses for the causal graph on ambiguous inputs.
Wrong outputs largely consist of copying, expected as the ambiguous examples are compatible with the Copying task. Note that, in the Parallel-only circuit, the activations at $y_{i}$ feeding into $y_{i}\rightarrow t_{N}$ are a mixture of activations for Copying or Present-Past; they each largely succeed in inducing the correct prediction on their own without contextualization when the task is consistent across the prompt, but $t_{N+1}$ fails at resolving conflicting task information when the task information is mixed across the prompt. We found that contextualization is needed to explain the model’s accuracy. Adding $x_{i}\rightarrow x_{i+1}$, $y_{i}\rightarrow y_{i+1}$ edges recovered part of the drop, but still underperformed the full model (68%). We next investigated which further edges help close the gap to the full model; indeed, in contrast to the standard tasks, we found separators now to play a role, as the full model with only separator-involving edges ablated performs at 85% (Table 9). Overall, adding a contextualization subcircuit with edges $y_{i}\rightarrow y_{j}$, $y_{i}\rightarrow t_{j}$, $x_{i}\rightarrow t_{j}$, $t_{i}\rightarrow y_{i}$, $t_{i}\rightarrow t_{i+1}$, $x_{i}\rightarrow x_{i+1}$ recovered almost all of the full model’s performance (89.9%). We investigated two hypotheses as to how contextualization might resolve ambiguity (Figure 5):
@@
We note that these hypotheses are not exclusive and might be true at the same time. To distinguish between them, we patched the K and V activations at either all ambiguous or all unambiguous examples feeding into $y_{i}\rightarrow t_{N+1}$ edges, with activations computed from either all-ambiguous or all-unambiguous prompts, in both cases erasing the presence of the other type of example (see Figure 23.2). Patching the $y_{i}$’s in ambiguous examples had little effect on the accuracy (Table 3, Figure 4). This suggests that, even if these examples may receive information about the presence of Present-Past examples, this is not causally important to task disambiguation. In contrast, patching the unambiguous examples had great effect: Accuracy dropped to almost zero when patching with ambiguous prompts (in which case, all task information would be routed through the ambiguous examples, but we saw that they do not causally provide disambiguation signal via $y_{i}\rightarrow t_{N+1}$). Strikingly, accuracy dropped drastically even when patching with fully unambiguous prompts, showing that contextualizing the unambiguous examples with awareness of the presence of ambiguous examples is key to downstream performance. Overall, our results allow us to reject H1. They are as expected under H2:
In the presence of ambiguous examples, contextualization changes the representations of unambiguous fewshot examples. These altered representations of unambiguous examples are key to transporting task information to $t_{N+1}$, overriding potentially misleading information from ambiguous examples.
@@ 4 Discussion
We have shown that Gemma-2 2B performs ICL using a cascade contextualizing the representations of individual examples, and aggregating task information from them. Using rigorous causal interventions, we extend prior work on mechanistically understanding the mechanisms of ICL in transformers (e.g. Wang et al., 2023; Cho et al., 2025; Kharlapenko et al., 2025; Hendel et al., 2023; Todd et al., 2024; Bansal et al., 2022; Song et al., 2025; Yin and Steinhardt, 2025; Crosbie and Shutova, 2024). We extend on work on function or task vectors by showing that they are largely part of the aggregation circuit hypothesized by Cho et al. (2025); Kharlapenko et al. (2025). Most importantly, we establish the importance of a contextualization step, especially in the presence of ambiguous examples in a prompt. A major difference between causal experiments in our work and Cho et al. (2025) is that we construct a circuit sufficient for recovering model performance, whereas Cho et al. (2025) zeroed out edges to confirm their importance. For instance, Cho et al. (2025) find that zeroing out edges involved in the aggregation circuit leads to a large drop in performance, but this does not show that these edges are also sufficient for achieving high performance. Our methodology allows us to obtain a faithful circuit recovering most of the model’s performance. A major difference to Kharlapenko et al. (2025) is that we keep different fewshot examples distinct in our circuits, allowing us to identify the contextualization step.
Theoretical work has considered ICL both from the perspective of learning simple parametric functions (e.g. Akyürek et al., 2023; Von Oswald et al., 2023; Mahankali et al., 2023), and as general Bayesian inference (e.g. Xie et al., 2022; Wies et al., 2023; Hahn and Goyal, 2023). Theoretical constructions for transformers performing ICL often correspond to the Parallel subcircuit (Akyürek et al., 2023; Mahankali et al., 2023), but contextualization of the fewshot inputs (Von Oswald et al., 2023; Vladymyrov et al., 2024) or outputs (Chen et al., 2024) does play a role in some constructions (see Appendix A for more), in agreement with our findings in an LLM. Mechanistic understanding of ICL also has the potential to explain some of its behavioral aspects. For instance, we found that, in the Person-Sport task, the circuit only assembles input and output spaces without regard to functional relationships, explaining the robustness to label noise sometimes observed (Min et al., 2022b).
@@ 5 Conclusion
We have conducted extensive causal interventions to understand the flow of information from individual fewshot examples to the next-token prediction in ICL, at the example of five tasks in Gemma-2 2B. Our most important result is to establish a two-stage procedure, whereby the model contextualizes representations of individual fewshot examples, which are then aggregated to prepare next-token prediction.

### v3

@@ 3.2 Which Information is Routed?
For each edge, we specifically contrast the following hypotheses: An edge from some few-shot example $(x_{i},y_{i})$ might transport causally relevant information about (1) the specific tokens, (2) the functional relationship between $x_{i}$ and $y_{i}$ (but not their token identities), (3) the type of the $x_{i}$ or $y_{i}$ tokens without regard to their functional relationship.
@@
We vary the information about the input that is available to a particular edge by constructing a contrastive input sample on which certain aspects of the information have been altered (Appendix H, Figure 22), and patching the K and V activations directly feeding into the edge with the corresponding activations computed on that contrastive input (Appendix, Figure 28). For instance, when patching an edge $y_{i}\rightarrow t_{N+1}$, we modify the key and value activations associated with the positions of $y_{i}$ across all attention heads in the model. These modifications occur specifically when the activations are queried from the positions corresponding to $t_{N+1}$. This patching is done in parallel in each layer and head. All patching is done within the overall circuit identified for the task, i.e., edges not present in the circuit are ablated, and paths not present in the overall circuit play no role even in the patched version. For instance, to determine whether the edges $y_{i}\rightarrow t_{N+1}$ transport contextualized information about preceding few-shots $(x_{i^{\prime}},y_{i^{\prime}})$ ($i^{\prime}<i$), we patch all inputs to this edge at position $y_{i}$ with activations computed on an input where the preceding few-shots have been corrupted; by varying the degree of corruption, we analyze which specific kinds of information are transported (Appendix H, Figure 22). We note that this methodology allows us to focus on causally relevant information used by downstream components. It is thus fundamentally different from supervised probing, which does not distinguish between information used or unused by downstream components. We focused on the 2B model for tractability.
Aggregation Subcircuit
@@
We first investigated the information transported by the edges in the Aggregation subcircuit, consisting of the $x_{i}\rightarrow y_{i}$, $x_{N+1}\rightarrow t_{N+1}$ and $y_{i}\rightarrow t_{N+1}$ edges. Prior work (Cho et al., 2025; Kharlapenko et al., 2025; Wang et al., 2023) suggests that the $y_{i}\rightarrow t_{N+1}$ edges aggregate task information from the examples. Indeed, when we patched these with arbitrary $x_{i}$ or $y_{i}$ (which disrupts the functional relationship), functionality was largely destroyed (see Figures 3 and 8 for all patching results). In contrast, when we patched the $y_{i}\rightarrow t_{N}$ edges with other few-shots that are valid for the task, functionality was preserved. We also verified that the $x_{i}\rightarrow y_{i}$ edges transport $x_{i}$: patching these edges with corrupted $x_{i}$ leads to total failure. This is to be expected, as in the simple Aggregation subcircuit, these edges are the only way for $x_{i}$ to influence the output. These results confirm prior proposals about the aggregation step (Cho et al., 2025; Kharlapenko et al., 2025; Wang et al., 2023) with rigorous causal analysis.
@@
Our methodology allowed us to obtain two further findings. First, the $y_{i}\rightarrow t_{N+1}$ edges causally transport only task information and no token information. Namely, even in the presence of contextualizing edges, i.e., even though there are multiple paths from few-shots to the $t_{N+1}$, performance did not deteriorate even when the $y_{i}\rightarrow t_{N+1}$ and contextualizing edges transported information from different prompts when they were valid for the task. [^9] We emphasize that this is a claim about information that has a causal downstream effect in the circuit, not about information that can be obtained via probing, which is likely to be richer and include token information.
@@
In order to understand the role of these components further, we investigated the functional behavior of the model with ablations applied. Overall, we observed three fallback tasks which accounted for most interpretable errors: copying the query, producing from the correct output space, copying another token from the prompt. With full ablation of the $y_{i}\rightarrow t_{N}$ edges, behavior was often accounted for by copying the query or copying another token from the corrupted input prompt (Appendix H, Figures 14, 15, 16, 17, 18). We note that as the corrupted prompts maintain the separators, the last separator position can still infer the presence of repeated structure, potentially explaining the presence of copying behavior. When patching the edge while preserving the output space, behavior often amounted to reproduction from the correct output space when there is well-delineated output space (e.g., Capitalization, Country-Capital, Person-Sport), or copying from the corrupted prompt when there is none (Copying). This is expected, as such prompts maintain repeated structure in terms of a specific output space, though with disrupted functional relationship.
@@
We next moved to analyzing the information transported by the contextualizing edges, starting with the edges connecting $y$’s ($y_{i}\rightarrow y_{i+1}$ or general $y_{i}\rightarrow y_{j}$). We found that these edges transport both information about $f$ and about $\mathcal{Y}$. Patching the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital only decreases accuracy significantly when $y$ leaves $\mathcal{Y}$, but not under other corruptions (Figure 3). In Capitalization, at 3 shots, replacing with $y\in\mathcal{Y}$ hurts much less than a general ablation of $y$ (accuracy 81% vs 56 %; Appendix F, Figure 8), which is as harmful as full corruption of these edges and leads to a large increase in copying behavior; nonetheless, accuracy still drops compared to ablations keeping $(x,y)\in f$. At 10 shots, Capitalization only responds to ablations where $y$ leaves $\mathcal{Y}$. In both cases, ablating the output space information leads to an increased number of copy-type mistakes (Appendix H, Figure 19).
@@
The $y_{i}\rightarrow y_{j}$ / $y_{i}\rightarrow y_{i+1}$ edges contextualize the representation of each few-shot example $y_{i}$ by transporting the type of $\mathcal{Y}$, and sometimes the task $f$, from $(x_{j},y_{j})$ ($j<i)$.
We next studied the edges connecting inputs $x$’s. Ablating these on Present-Past and Person-Sport turns out to not lead to a statistically significant drop, but the drop is substantive in Country-Capital (Figure 3). Here, ablations only hurt when $x$ left $\mathcal{X}$, showing that $x_{i}\rightarrow x_{j}$ transports the input space, but not (causally) the token $x_{i}$. Such ablations again lead to a substantial fraction of copying responses, i.e., the functional input-output behavior is impacted (Appendix H, Figure 20). Patching prior $x_{i}$’s for these edges, we found that they affect the final prediction both via $x_{i}\rightarrow x_{i+1}\rightarrow y_{i+1}\rightarrow t_{N+1}$ and via $x_{N}\rightarrow x_{N+1}\rightarrow t_{N+1}$ edges (Appendix, Table 8). Overall:
The $x_{i}\rightarrow x_{i+1}$ edges contextualize the representations of $x_{N+1}$ and of individual few-shot outputs $y_{i}$ with information about $\mathcal{X}$.
Figure 4: Two hypotheses for the causal graph on ambiguous inputs.
We conducted a controlled experiment aiming to stress-test the aggregation step. We designed the following hard version of the Present-Past task. We take advantage of the fact that some English verbs have identical present and past tense forms (e.g., “put”, “spread”). In a prompt for the Present-Past task, examples using such verbs are thus ambiguous between the Present-Past task and a simpler Copying task (Appendix H, Figure 23.1). In this experiment, most individual examples do not give full task information, and may in fact provide misleading information. We hypothesize that direct aggregation of uncontextualized per few-shot representations as hypothesized by prior work may be less useful in this case, because individual examples might, on their own, provide conflicting task information. The contextualization step might help resolve this ambiguity before the aggregation step.
Even when 7 out of 10 few-shot examples are ambiguous, the full model continues to perform the task at very high accuracy (95%). Whereas the aggregation (Aggregation)-only circuit had performed at 93% on the standard Present-Past task (and similarly on the Copying task), its performance on the ambiguous prompts dropped to 56% (Appendix H, Table 9).
@@ 3.3 A Contextualization subcircuit is important in the presence of ambiguity
Wrong outputs largely consist of copying, expected as the ambiguous examples are compatible with the Copying task. Note that, in the Aggregation-only circuit, the activations at $y_{i}$ feeding into $y_{i}\rightarrow t_{N}$ are a mixture of activations for Copying or Present-Past; they each largely succeed in inducing the correct prediction on their own without contextualization when the task is consistent across the prompt, but $t_{N+1}$ fails at resolving conflicting task information when the task information is mixed across the prompt.
We found that contextualization is needed to explain the model’s accuracy. Adding $x_{i}\rightarrow x_{i+1}$, $y_{i}\rightarrow y_{i+1}$ edges recovered part of the drop, but still underperformed the full model (68%). We next investigated which further edges help close the gap to the full model; indeed, in contrast to the standard tasks, we found separators now to play a role, as the full model with only separator-involving edges ablated performs at 85% (Appendix H, Table 9). Overall, adding a contextualization subcircuit with edges $y_{i}\rightarrow y_{j}$, $y_{i}\rightarrow t_{j}$, $x_{i}\rightarrow t_{j}$, $t_{i}\rightarrow y_{i}$, $t_{i}\rightarrow t_{i+1}$, $x_{i}\rightarrow x_{i+1}$ recovered almost all of the full model’s performance (89.9%).
We investigated two hypotheses as to how contextualization might resolve ambiguity (Figure 4):
@@
We note that these hypotheses are not exclusive and might be true at the same time. To distinguish between them, we patched the K and V activations at either all ambiguous or all unambiguous examples feeding into $y_{i}\rightarrow t_{N+1}$ edges, with activations computed from either all-ambiguous or all-unambiguous prompts, in both cases erasing the presence of the other type of example (see Appendix H, Figure 23.2). Patching the $y_{i}$’s in ambiguous examples had little effect on the accuracy (Table 3, Figure 5). This suggests that, even if these examples may receive information about the presence of Present-Past examples, this is not causally important to task disambiguation.
In contrast, patching the unambiguous examples had great effect: Accuracy dropped to almost zero when patching with ambiguous prompts (in which case, all task information would be routed through the ambiguous examples, but we saw that they do not causally provide disambiguation signal via $y_{i}\rightarrow t_{N+1}$). Strikingly, accuracy dropped drastically even when patching with fully unambiguous prompts, showing that contextualizing the unambiguous examples with awareness of the presence of ambiguous examples is key to downstream performance. Overall, our results allow us to reject H1. They are as expected under H2:
In the presence of ambiguous examples, contextualization changes the representations of unambiguous few-shot examples. These altered representations of unambiguous examples are key to transporting task information to $t_{N+1}$, overriding potentially misleading information from ambiguous examples.
@@ 4 Discussion
On five tasks, we have shown that Gemma-2 2B performs ICL using a cascade contextualizing the representations of individual examples, and aggregating task information from them. Using rigorous causal interventions, we extend prior work on mechanistically understanding the mechanisms of ICL in transformers (e.g. Wang et al., 2023; Cho et al., 2025; Kharlapenko et al., 2025; Hendel et al., 2023; Todd et al., 2024; Bansal et al., 2022; Sia et al., 2024; Song et al., 2025; Yin and Steinhardt, 2025; Crosbie and Shutova, 2024). We extend on work on function or task vectors by showing that they are largely part of the aggregation circuit hypothesized by Cho et al. (2025); Kharlapenko et al. (2025). Most importantly, we establish the importance of a contextualization step, especially in the presence of ambiguous examples in a prompt.
Our work is distinct from Cho et al. (2025) in providing a computation graph supported by causal effects rather than representation similarity. A major difference between causal experiments in our work and Cho et al. (2025) is that we construct a circuit sufficient for recovering model performance, whereas Cho et al. (2025) zeroed out edges to confirm their importance. For instance, Cho et al. (2025) find that zeroing out edges involved in the aggregation circuit leads to a large drop in performance, but this does not show that these edges are also sufficient for achieving high performance. Our methodology allows us to obtain a faithful circuit recovering most of the model’s performance. A major difference to Kharlapenko et al. (2025) is that we keep different few-shot examples distinct in our circuits, allowing us to identify the contextualization step.
Theoretical work has considered ICL both from the perspective of learning simple parametric functions (e.g. Akyürek et al., 2023; Von Oswald et al., 2023; Mahankali et al., 2023), and as general Bayesian inference (e.g. Xie et al., 2022; Wies et al., 2023; Hahn and Goyal, 2023). Theoretical constructions for transformers performing ICL often correspond to the Aggregation subcircuit (Akyürek et al., 2023; Mahankali et al., 2023), but contextualization of the few-shot inputs (Von Oswald et al., 2023; Vladymyrov et al., 2024) or outputs (Chen et al., 2024) does play a role in some constructions (see Appendix A for more), in agreement with our findings in an LLM. Mechanistic understanding of ICL also has the potential to explain some of its behavioral aspects. For instance, we found that, in the Person-Sport task, the circuit only assembles input and output spaces without regard to functional relationships, explaining the robustness to label noise sometimes observed (Min et al., 2022).
@@ 5 Conclusion
We have conducted extensive causal interventions to understand the flow of information from individual few-shot examples to the next-token prediction in ICL, at the example of five tasks in Gemma-2 2B. Our most important result is to establish a two-stage procedure, whereby the model contextualizes representations of individual few-shot examples, which are then aggregated to prepare next-token prediction.